40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 57 of 330  Not logged in ELOG logo
ID Date Author Type Category Subject
  13784   Tue Apr 24 11:31:59 2018 gautamConfigurationALSProposed changes to EX fiber coupling

Motivation: I want to make another measurement of the out-of-loop ALS beat noise, with improved MM into both the PSL and EX fibers and also better polarization control. For this, I want to make a few changes at the EX table. 

  1. Replace existing fiber collimator with one of the recently acquired F220-APC-1064 collimators.
    • This gives an output mode of diameter 2.4mm with a beam divergence angle of 0.032 degrees (all numbers theoretical - I will measure these eventually but we need a beam path of ~5m length in order to get a good measurement of this collimated beam).
    • I believe it will be easier to achieve good mode matching into this mode rather than with the existing collimator. 
    • Unfortunately, the mount is still going to be K6X and not K6XS. 
  2. Improve mode-matching into fiber.
    • I used my measurement of the Innolight NPRO mode from 2016, a list of available lenses, and some measured distances to calculate a solution that gives theoretical 100% overlap with the collimator mode, that has beam diameter 2.4mm, located 80cm from the NPRO shutter head location (see Attachment #1).
    • The required movement of components is schematically illustrated in Attachment #2.
    • One of the required lens positions is close to the bracket holding the enclosure to the table, but I think the solution is still workable (the table is pretty crowded so I didn't bother too much with trying to find alternative solutions as all of them are likely to require optics placed close to existing ones and I'd like to avoid messing with the main green beam paths.
    • I will attempt to implement this and see how much mode matching we actually end up getting.
  3. Install a PBS + HWP combo in the EX fiber coupling path.
    • This is for better polarization control.
    • Also gives us more control over how much light is coupled into the fiber in a better way than with the ND filters in current path.
  4. Clean EX fiber tip.
  5. Dump a leakage IR beam from the harmonic separator post doubling oven, which is currently just hitting the enclosure. It looks pretty low power but I didn't measure it.
  6. Re-install EX power monitoring PD.

Barring objections, I will start working on these changes later today.

Attachment 1: EX_fiber_MM.pdf
EX_fiber_MM.pdf
Attachment 2: EX_fiber_changes.png
EX_fiber_changes.png
  13783   Tue Apr 24 10:10:43 2018 gautamUpdateComputer Scripts / ProgramsParticle swarm hyper parameter optimization

I'm copying and pasting Nikhil's email here as he was unable to login to the elog (but should now be able to in order to reply to any comments, and add more details about this test, motivation, methodology etc).

I did some post-processing after running the grid search. The following steps were carried out:

1) Selected those sets whose cost fun were less than a specific threshold (here 10000)

2) Next task was to see if the parameters of these good solutions had some pattern

3) I used a dimensionality reduction technique called t-SNE to project the 6 dimensional parameter space to 2 dim (for better visualization )

4) Made a scatter plot of these (see fig )

5) Used K-Means to find the clusters in this data

6) MarkerSize & Color reflect the cost fun. Bigger the marker size means better the solution.

7) Visual inspection implied cluster 5 had the best ranking points & more than any other cluster

8) These points had the following Parameter set: Workers {20,40}, SwarmSize {500}, MaxIter {500}, Self Adjustment {1}, Social Adjustment {1}, Tolerance {1e-3,1e-8} 

     See fig: for the box plot 

9) It looks like is a particular set of values rather than individual values that gives the best results.

 

Attachment 1: ClusterFminScaled.png
ClusterFminScaled.png
Attachment 2: ClusterID_5.png
ClusterID_5.png
  13782   Tue Apr 24 09:10:20 2018 KiraUpdatePEMfinal setup sketch

I've attached the final sketch for the panel on the granite block.

Attachment 1: EX-can-panel_1.pdf
EX-can-panel_1.pdf
  13781   Tue Apr 24 08:36:47 2018 johannesConfigurationGeneralAux Laser LD dying? (AS port laser injection)

In September 2017 I measured ~150mW output power, which was already kind of low. What are the chances of getting this one repaired? Steve, can you please check the serial number? It's probably too old like the other ones.

Quote:

I suspect that the LD of the aux laser is dying.
- The max power we obtain from this laser (700mW NPRO) is 33mW. Yes, 33mW. (See attachment 1)

 

  13780   Mon Apr 23 20:06:35 2018 ranaUpdatePEMPID tune

This shows a step response of the EX seis temp control with K_I = -1 and K_P = -0.1. The time constants for both heatup and cooldown are ~2 hours.

I'm not so sure if the PID code itself makes sense though:

  # The basic finite-difference PID approximation
  e[0] = (p-s)
  print("Error signal = {}" .format(e[0])) 

  # These are the main equations of the PID Process
  u[0] = u[1]
  u[0] = u[0] + Kp * (e[0] - e[1])
  u[0] = u[0] + Ki * (e[0])
  u[0] = u[0] + Kd * (e[0] - 2*e[1] + e[2])

 

Seems like the Proportional term uses the difference (or derivative) of the error signal. This makes it more likely to pick up some high frequency noise; maybe we should low pass this signal somewhat, or at least implement a running average.

Since we still don't have an out of loop sensor or a PSL room temperature monitor or a particle counter in the frames, I've disabled the PID loop to see how much the can temperature varies with no feedback. Please leave it this way for a few days.

Attachment 1: HeaterTest.png
HeaterTest.png
  13779   Sat Apr 21 20:25:12 2018 gautamUpdateALSPSL fiber pickoff status

Seems like there is still a bit of variation in the power in the two polarizations, though it is much smaller now, at the ~5% level (see Attachment #1). Since the pattern is repeating itself over the day timescale, I think this effect is not because of rotation of the output coupler in the mount, but is in fact a temperature driven waveplate effect because of imperfect alignment at the input coupler (which itself is locked down). I'm going to rotate the input coupler by 5 degrees (old = 110 degrees, new=115degrees) to see if the situation improves...


gautam Apr 24 2pm: Steve suggested confirming the correlation by hooking up the PSL table temperature sensor. This used to be logged but since the c1psl ADC card failure, has not been recorded. Assuming the sensor and preamp still work fine, we can use the PSL diagnostic Acromag (whose channels I have hijacked to monitor polarization stability already) to at least temporarily monitor the temperature inside the PSL enclosure. I am in need of a DB25 breakout board for this purpose which I am missing right now, as soon as I obtain one, I'll hook this up...

Attachment 1: PSL-beatMouthPickoff.png
PSL-beatMouthPickoff.png
  13778   Sat Apr 21 20:19:05 2018 gautamUpdateGeneralMegatron hard-rebooted

I found megatron in a similar state to that which nodus was in yesterday. Clued by the fact that MCautolocker wasn't executing the mc scripts (as was evident from looking at the wall StripTool trace), I tried ssh-ing into megatron, but was unable to (despite it being responsive to ping requests). So I went into the VEA and plugged in a monitor to megatron - saw nothing on it. With no soft reboot options available, I power cycled the machine via the front panel button. It came back up smoothly. I manually restarted the autolocker, FSSslow and EX thermal control processes (the former two with initctl, while the latter runs in a tmux session). Everything seems alright for now. Not sure how long megatron has been dead for.

  13777   Fri Apr 20 23:36:28 2018 KevinUpdatePEMSeismometer BLRMs

Steve secured the GPS time server in the rack above the AA board and removed the wooden block that it was resting on. The new rack is shown in attachment 1.

I then opened the AA board to see why the channels aren't working. Even though the board was powered and outputting 4.6 V, none of the chips were getting power. I must have shorted something while trying to diagnose this and the board is no longer powered either.

The schematic is given in D990147. The D68L8EX filter is bypassed on all the channels, as can be seen in attachment 3, so the board isn't really doing anything. Rana suggested that we could just bypass the whole circuit by wiring the IN channels directly to the OUT channels going to the ADC. I'll try that next for a single channel.

Attachment 1: front.jpg
front.jpg
Attachment 2: back.jpg
back.jpg
Attachment 3: detail.jpg
detail.jpg
  13776   Fri Apr 20 16:25:08 2018 SteveUpdateWiki ETMX table layout uploaded to wiki

ETMX table layout uploaded with beam paths to the wiki.   laugh

  13775   Fri Apr 20 16:22:32 2018 gautamUpdateGeneralNodus hard-rebooted

Aidan called saying nodus was down at ~345pm. I was able to access it at ~330pm. I couldn't ssh in from my machine or the control room ones. So I went to 1X7 and plugged in a monitor to nodus. It was totally unresponsive. Since the machine wasn't responding to ping either, I decided to hard-reboot it. Machine seemed to come back up smoothly. I had trouble getting the elog started - it wasn't clear to me that the web ports were closed by default, so even though the startELOGD.sh script was running fine, the 8080 port wasn't open to the outside world. Anyways, once I figured this out, I was able to start the elog. DokuWiki also seems to be up and running now... 

  13774   Fri Apr 20 15:07:45 2018 KiraUpdatePEMfinal setup sketch update

If we lay the cable along the floor then it should be around 6' to the current setup and about 20' to the actual seismometer.

Edit: 16 gauge wire should be good.

Quote:

since we're just going from the short rack (not the tall rack) to the seismometer, can't we use a cable shorter than 45' ?

 

  13773   Fri Apr 20 00:26:34 2018 gautamUpdateALSFibers switched out

Summary:

I think the dominant cause for the fact that we were seeing huge swing in the power coupled into the fiber was that the beam being sent in was in fact not linearly polarized, but elliptically polarized. I've rectified this with the help of a PBS. Fiber has been plugged into my polarization monitoring setup. Let's monitor for some long stretch and see if the situation has improved.

Details:

  • The new fiber mount I ordered, K6XS, arrived today. I like it - it has little keys with which all DoFs can be locked. Moreover, it is compatible with the fixed collimators which IMO is the easiest way to achieve good mode-matching into the fiber. It is basically a plug-and-play replacement for the mounts we were using. Anyways, we can evaluate the performance over the coming days.
  • I installed it on the PSL table (started work ~10pm, HEPA turned up to maximum, PSL shutter closed).
  • But even with the new rotational DoF locking feature, I saw that slight disturbances in the fiber caused wild fluctuations in my polarization monitoring setup PD outputs. This was a useful tool through the night of checking the polarization content in the two special axes - Aidan had suggested using a heat gun but shaking the fiber a bit works well too I think.
  • The PM980 fiber has an alignment key that is aligned with the slow axis of the fiber - so it is a useful alignment reference. But even by perturbing the roational alignment about the vertical by +/-15 degrees, I saw no improvement in this behavior. So I began to question my assumption that the input beam itself had clean polarization content.
  • Since my pickoff beam has gone through a QWP and two PBSs, I had assumed that the beam was linearly polarized.
  • But by putting a PBS just upstream of the input fiber coupler, I could see a beam at the S-port with an IR card (while I expected the beam to be P-polarized).
  • OK - so I decided to clean up the input polarization by leaving this PBS installed. With this modification to the setup, I found that me shaking the fiber around on the PSL table didn't affect the output polarization content nearly as dramatically as before!!yes
  • The state I am leaving it in tonight is such that there is ~100x the power in the P-polarization output monitor as the S-polarization (PER ~ 20dB). I didn't try and optimize this too much more for now, I want to observe some long term trend to see if the wild power fluctuations have been mitigated.
  • The output coupler is mounted on the inferior K6X mount, and so there is the possibility that some drift will be attributable to rotation of the output coupler in it's mount. Thermally driven length changes / time varying stresses in the fiber may also lead to some residual power fluctuations. But I don't expect this to be anywhere near the ~25% I reported in the previous elog.
  • The rejected beam from the PBS was measured to be ~300 uW. I haven't dumped this properly, to be done tomorrow.
  • HEPA turned back down to 30%, PSL enclosure closed up, PSL shutter re-opened ~0030am.
  • Note that the EX and EY fiber coupled beams are also likely subject to the same problem. We have to double check. I think it's better to have a PBS in front of the input fiber coupler as this also gives us control over the amount of light coupled into the fiber.

Power budget:

Power in Measured power (Ophir, filter OFF)
@Input coupler, before PBS 4.4 mW
P-pol content @ input coupler 4.06 mW
S-pol (rejected) from PBS 275 uW
@Output coupler 2.6 mW (MM ~65%)

 

  13772   Thu Apr 19 20:41:09 2018 KojiConfigurationGeneralAux Laser LD dying? (AS port laser injection)

I suspect that the LD of the aux laser is dying.
- The max power we obtain from this laser (700mW NPRO) is 33mW. Yes, 33mW. (See attachment 1)
- The intensity noise is likely to be relaxation oscillation and the frequency is so low as the pump power is low. When the ADJ is adjusted to 0, the peak moved even lower. (Attachment 2, compare purple and red)
- What the NE (noise eater) doing? Almost nothing. I suspect the ISS gain is too low because of the low output power. (Attachment 2, compare green and red)

Attachment 1: Aux_laser_adj_Pout.pdf
Aux_laser_adj_Pout.pdf
Attachment 2: Aux_laser_RIN.pdf
Aux_laser_RIN.pdf
  13771   Thu Apr 19 18:23:51 2018 KiraUpdatePEMfinal setup sketch update

since we're just going from the short rack (not the tall rack) to the seismometer, can't we use a cable shorter than 45' ?

Quote:

I've added in the dimensions to my sketch.

the panel should be completely replaced like I described. We don't want to try to squeeze it in artificially and torque the wires. It just needs to be separated from the slab by a few more cm.

  13770   Thu Apr 19 17:15:35 2018 gautamUpdateIOOMore IMC NBing

Summary:

Today, I repeated the coil driver noise measurement. Now, the coil driver noise curve in the noise budget plot (Attachment #1) is the actual measurement of all 12 coils (made with G=100 SR560). I am also attaching the raw voltage noise measurement (input terminated in 50ohms, Attachment #2). Note that POX11 spectrum has now been re-measured with analog de-whitening engaged on all 3 optics such that the DAC noise contribution should be negligible compared to coil driver noise in this configuration. The rows in Attachment #2 correspond to 800 Hz span (top) and full span (bottom) on the FFT analyzer.

Details:

The main difference between this measurement, and the one I did middle of last year (which agreed with the expectation from LISO modeling quite well) is that this measurement was done in-situ inside the eurocrate box while last year, I did everything on the electronics benches. So I claim that the whole mess of harmonics seen in the raw measurements are because of some electronics pickup near 1X6. But even disregarding the peaky features, the floor of ~30nV/rtHz is ~6x than what one would expect from LISO modeling (~5nV/rtHz). I confirmed by looking that the series resistance for all 3 MC optics is 430ohms. I also did the measurement with the nominal bias voltages applied to the four channels (these come in via the slow ADCs). But these paths are low-passed by an 8th order low pass with corner @ 1Hz, so at 100 Hz, even 1uV/rtHz should be totally insignificant. I suppose I could measure (with EPICS sine waves) this low-pass filtering, but it's hard to imagine this being the problem. At the very least, I think we should get rid of the x3 gain on the MC2 coil driver de-whitening board (and also on MC1 and MC3 if they also have the x3 factor).

Attachment 1: IMC_NB_20180419.pdf
IMC_NB_20180419.pdf
Attachment 2: IMC_coilDriverNoises_20180419.pdf
IMC_coilDriverNoises_20180419.pdf
  13769   Thu Apr 19 12:23:30 2018 KiraUpdatePEMfinal setup sketch update

I've added in the dimensions to my sketch.

It seems like placing the two connectors right next to each other would allow both cables to just barely go through the hole in the block.

Quote:

Can you please add dimensions to the drawing, so we can see if things fit and what the cable lenghts need to be?

For the panel on the granite slab, we should use a thinner piece of metal and mount it with an offset so that the D-sub cable can be fished through the hole in the slab. The hole is wide enough for 2 cables, but not 2 connectors.

 

Attachment 1: heater_1_new.png
heater_1_new.png
Attachment 2: heater_2_new.png
heater_2_new.png
  13768   Thu Apr 19 11:29:11 2018 ranaUpdatePEMPID tune

Yesterday, I changed the P gain of the PID loop from zero to  +0.1. Seems good so far; will monitor for a couple days to see if we're in the right ballpark. Main issue in the stability may now be that the quantization noise is too big for the temperature sensor. If so, we should consider subtracting off the DC value (with a V ref) and then amplifying before ADC.

Attachment 1: Screen_Shot_2018-04-19_at_11.27.08_AM.png
Screen_Shot_2018-04-19_at_11.27.08_AM.png
  13767   Thu Apr 19 09:57:03 2018 gautamUpdateWikiAP and ETMX tables uploaded to wiki

The most up to date pictures of the AP table and ETMX table that Steve took have been uploaded to the relevant page on the wiki. It seems like the wiki doesn't display previews of jpg images - by using png, I was able to get the thumbnail of the attachment to show up. It would be nice to add beam paths to these two images. The older versions of these photos were moved to the archive section on the same page.

  13766   Thu Apr 19 01:04:00 2018 gautamConfigurationGeneralAS port laser injection

For the arm cavity ringdowns, I guess we don't need AS55/AS110 (although I think the camera will still be useful for alignment). But for something like RC Gouy phase characterization, I'd imagine we need the AS detectors to lock various cavities. So I think we should go for a solution that doesn't disturb the AS PD beams. 

It's hard to tell from the plot in the manual (pg 52) what exactly the relaxation oscillation frequency is, but I think it's closer to 600 kHz (is this characteristic of NdYAG NPROs)??  Is the high RIN on the light straight out of the NPRO? 

Quote:

We could use a simple mirror if we don't need AS110 and AS55, we could use a polarizing BS and work with s polarization, or we find a Faraday Isolator.


There is a strong oscillation around 210 kHz. The oscillation frequency decreases when the output power is turned down, the noise eater has no effect. 

  13765   Thu Apr 19 00:03:51 2018 gautamUpdateIOOMore IMC NBing

Summary:

As shown in the Attachments, it seems like IMC DAC and coil driver noise is the dominant noise source above 30Hz. If we assume the region around the bounce peak is real motion of the stack (to be confirmed with accelerometer data soon), this NB is starting to add up. Much checking to be done, and I'd also like to get a cleaner measurement of coil driver and DAC noise for all 3 optics, as there seems to be a factor of ~5 disagreement between the MC3 coil driver noise measurement and my previous foray into this subject. The measurement needs to be refined a little, but I think the conclusion holds.

Details:

  1. I had a measurement of the MC3 coil driver noise from ~2weeks ago when I was last working on this that I had not yet added to the NB.
  2. Today I added it. To convert from measured voltage noise to frequency noise, I assumed the usual 0.016N/A per coil number, which is probably a large source of systematic error.
  3. I define the "nominal" IMC operating condition as MC1 and MC3 having the analog de-whitening filters switched on, but MC2 switched off.
  4. So length noise should be dominated by coil driver noise on MC1 and MC3, and DAC noise on MC2.
  5. The measurement I had was made with the input to the coil driver board terminated in 50ohms. Measurement was made in-situ. The measurement has a whole bunch of 60Hz harmonics (despite the Prologix box being powered by a linear power adapter, but perhaps there are other ground loops which are coupling into the measurement). So I'd like to get a cleaner measurement tmrw.
  6. To confirm, Koji suggested some On/Off test by driving some broadband noise in the coils. I figured toggling the analog de-whitening, such that the DAC noise or coil driver electronics dominate is an equally good test.
  7. Attachment #2 shows the effect in arm error and control signal spectra. Note that I engaged analog de-whitening on all 3 optics for the red curves in this plot. But even leaving MC2 de-whitening off, I could see the read curve was below the black reference trace, which was taken with de-whitening off on all 3 optics.

Remarks:

Since I sunk some time into it already, the motivation behind this work is just to try and make the IMC noise budget add up. It is not directly related to lowering the IR ALS noise, but if it is true that we are dominated by coil driver noise, we may want to consider modifying the MC coil driver electronics along with the ITM and ETMs.

Attachment 1: IMC_NB_20180409.pdf
IMC_NB_20180409.pdf
Attachment 2: IMC_coils_20180418.pdf
IMC_coils_20180418.pdf
  13764   Wed Apr 18 22:46:23 2018 johannesConfigurationGeneralAS port laser injection

Using Gautam's Finesse file and the cad files for the 40m optical setup I propagated the arm mode out of the AS port. For the location of the 3.04 mm waist I used the average distance to the ITMs, which is 11.321 m from the beam spot on the 2 inch mirror on the AS table close to the viewport. The 2inch lens focuses the IFO mode to a 82.6 μm waist at a distance of 81 cm, which is what we have to match the aux laser fiber output to.

I profiled the fiber output and obtained a waist of 289.4 μm at a distance of 93.3 cm from the front edge of the base of the fiber mount. Next step is to figure out the lens placement and how to merge the beam paths. We could use a simple mirror if we don't need AS110 and AS55, we could use a polarizing BS and work with s polarization, or we find a Faraday Isolator.


While doing a beam scan with the razor blade method I noticed that the aux laser has significant intensity noise. This is seen on the New Focus 1611 that is used for the beat signal between PSL and aux laser, as well as on the fiber output PD. There is a strong oscillation around 210 kHz. The oscillation frequency decreases when the output power is turned down, the noise eater has no effect. Koji suggested it could be light scattering back into the laser because I couldn't find a usable Faraday Isolator back when I installed the aux laser in the PSL enclosure. I'll have to investigate this a little further, look at the spectrum, etc. This intensity noise will appear as amplitude noise of the beat note, which worries me a little.

power_out_fluctuation_DC.png      power_out_fluctuation_AC_zoom.png

Attachment 1: ASpath.svg.png
ASpath.svg.png
  13763   Wed Apr 18 20:33:19 2018 KevinUpdateGeneralseismometer interfaces

Steve, the pictures you posted are not the AA board I was referring to. The attached pictures show the board which is sitting beneath the GPS time server.

Attachment 1: front.jpg
front.jpg
Attachment 2: back.jpg
back.jpg
Attachment 3: connectors.jpg
connectors.jpg
  13762   Wed Apr 18 19:02:15 2018 gautamUpdateIOOMC spot centering scripts

I'm working on fixing these (and the associated MEDM scripts) up so that we can get some reliable readback on how well centered the spots are on the MC mirrors. Seems like a bunch of MEDM display paths were broken, and it looks like the optimal demod phases (to put most of the output in I quadrature) are not what the current iteration of the scripts set them to be. It may well be that the gains that convert demodulated counts to mm of spot offset are also not correct anymore. I ran the script ~4times in ~1 hour time span, and got wildly different answers for the spot centering each time, so I wouldn't trust any of those numbers atm...

As you can see in Attachment #1, I stepped the demod phase of one of the servos from -180 to 180 degrees in 5degree increments. The previously used value of 57degrees is actually close to the worst possible point (if you want the signal in the I quadrature, which is what the scripts assume).

I used Attachment #2 to change up the demod phases to maximize the I signal. I chose the demod phase such that it preserved the sign of the demodulated signal (relative to the old demod phases). I also made some StripTool templates for these, and they are in the MC directory. Doing the spot centering measurement with the updated demod phases yields the following output from the script:

spot positions in mm (MC1,2,3 pit MC1,2,3 yaw):
[0.72506586251144134, 7.1956908534034403, 0.54754354165324848, -0.94113388241389784, -3.5325883025464959, -2.4934027726657142]

Seems totally unbelievable still that we are so far off center on MC1 yaw. Perhaps the gains and calibration to convert from counts to mm of spot offset need to be rechecked.

Attachment 1: MCass_demodPhase.png
MCass_demodPhase.png
Attachment 2: MCass_demodPhases.png
MCass_demodPhases.png
  13761   Wed Apr 18 17:15:35 2018 ranaConfigurationComputersNODUS: no xmgrace for dataviewer

Turns out, there is no RPM for XmGrace on Scientific Linux 7. Since this is the graphic output of dataviewer, we can't use dataviewer through X windows until this gets fixed. CDS is looking into a xmGrace replacement, but it would be better if we can hijack a alt RH repo to steal a temporary xmgrace RPM. KT has been pinged.

  13760   Wed Apr 18 16:59:35 2018 ranaUpdatePEMfinal setup sketch: EX Seis

Can you please add dimensions to the drawing, so we can see if things fit and what the cable lenghts need to be?

For the panel on the granite slab, we should use a thinner piece of metal and mount it with an offset so that the D-sub cable can be fished through the hole in the slab. The hole is wide enough for 2 cables, but not 2 connectors.


Attached is a 8-day minute trend of the heater control signals, as well as the in-loop temperature sensor (which underestimates the true fluctuations; we really need an out-of-loop sensor attached to the can or seismometer).

You can see that since the last tuning (on the 13th), its been stable at the set point of 39 C with 8.5 - 10 W of heating power. Need to add the PID loop settings (all the sliders on the MEDM screen) to the frames so that we can help in diagnosing. Also, fix the spelling of "Celcisususs".

Attachment 1: Screen_Shot_2018-04-18_at_5.20.53_PM.png
Screen_Shot_2018-04-18_at_5.20.53_PM.png
  13759   Wed Apr 18 12:18:39 2018 KiraUpdatePEMfinal setup sketch

I've updated the sketches and added in front panels for the seismometer block and the 1U panel (attachments 3 and 4). There was an issue when it came to the panel on the block because the hole is only big enough for the cable that already exists there and there is no space to add in the D-9 connector. Not quite sure how to resolve this issue. Attachment 7 is the current panel on the seismometer block. Attachments 5 and 6 are the updated temperature circuit and the heater circuit.

The boxes will be located in the short racks at EX and EY to minimize cable length.

Attachment 1: heater_1_new.png
heater_1_new.png
Attachment 2: heater_2_new.png
heater_2_new.png
Attachment 3: 1U-panel.pdf
1U-panel.pdf
Attachment 4: EX-can-panel.pdf
EX-can-panel.pdf
Attachment 5: IMG_20180412_120427.jpg
IMG_20180412_120427.jpg
Attachment 6: HeaterCircuit.pdf
HeaterCircuit.pdf
Attachment 7: IMG_20180418_121115.jpg
IMG_20180418_121115.jpg
  13758   Wed Apr 18 10:44:45 2018 gautamUpdateCDSslow machine bootfest

All slow machines (except c1auxex) were dead today, so I had to key them all. While I was at it, I also decided to update MC autolocker screen. Kira pointed out that I needed to change the EPCIS input type (in the RTCDS model) to a "binary input", as opposed to an "analog input", which I did. Model recompilation and restart went smooth. I had to go into the epics record manually to change the two choices to "ENABLE" and "DISABLE" as opposed to the default "ON" and "OFF". Anyways, long story short, MC autolocker controls are a bit more intuitive now I think.

Attachment 1: MCautolocker_MEDM_revamp.png
MCautolocker_MEDM_revamp.png
  13757   Tue Apr 17 14:08:29 2018 gautamUpdateALSFibers switched out

A follow-up on the discussion from today's lunch meeting - Rana pointed out that rotation of the fiber in the mount by ~5degrees cannot account for such large power fluctuations. Here is a 3 day trend from my polarization monitoring setup. Assuming the output fiber coupler rotates in its mount by 5 degrees, and assuming the input light is aligned to one of the fiber's special axes, then we expect <1% fluctuation in the power. But the attached trend shows much more drastic variations, more like 25% in the p-polarization (which is what I assume we use for the beat, since the majority of light is in this polarization, both for PSL and EX). I want to say that the periodicity in the power fluctuations is ~12hours, and so this fluctuation is somehow being modulated by the lab temperature, but unfortunately, we don't have the PSL enclosure temperature logged in order to compare coherence.

Steve: your  plots look like temperature driven


The "beat length" of the fiber is quoted as <=2.7mm. This means that a linearly polarized beam that is not oriented along one of the special axes of the fiber will be rotated through 180 degrees over 2.7mm of propagation through the fiber. I can't find a number for the coefficient of thermal expansion of the fiber, but if temperature driven fluctuations are changing the length of the fiber by 300um, it would account for ~12% power fluctuation between the two polarizations in the monitoring setup, which is in the ballpark we are seeing...

Attachment 1: PSL_fluctuations.png
PSL_fluctuations.png
  13756   Tue Apr 17 09:57:09 2018 SteveUpdateGeneralseismometer interfaces

 

Quote:

I've been looking into recovering the seismic BLRMs for the BS Trillium seismometer. It looks like the problem is probably in the anti-aliasing board. There's some heavy stuff sitting on top of it in the rack, so I'll take a look at it later when someone can give me a hand getting it out.

In detail, after verifying that there are signals coming directly out of the seismometer, I tried to inject a signal into the AA board and see it appear in one of the seismometer channels.

  1. I looked specifically at C1:PEM-SEIS_BS_Z_IN1 (Ch9), C1:PEM-SEIS_BS_X_IN1 (Ch7), and C1:PEM-ACC_MC2_Y_IN1 (Ch27). All of these channels have between 2000--3000 cts.
  2. I tried injecting a 200 mVpp signal at 1.7862 Hz into each of these channels, but the the output did not change.
  3. All channels have 0 cts when the power to the AA board is off.
  4. I then tried to inject the same signal into the AA board and see it at the output. The setup is shown in the first attachment. The second BNC coming out of the function generator is going to one of the AA board inputs; the 32 pin cable is coming directly from the output. All channels give 4.6 V when when the board is powered on regardless of wheter any signal is being injected.
  5. To verify that the AA board is likely the culprit, I also injected the same signals directly into the ADC. The setup is shown in the second attachment. The 32 pin cable is going directly to the ADC. When injecting the same signals into the appropriate channels the above channels show between 200--300 cts, and 0 cts when no signal is injected.

 

Attachment 1: BS_Tril_Intrf-1X5.jpg
BS_Tril_Intrf-1X5.jpg
Attachment 2: Gurs_Intf-1X1.jpg
Gurs_Intf-1X1.jpg
  13755   Mon Apr 16 22:09:53 2018 KevinUpdateGeneralpower outage - BLRM recovery

I've been looking into recovering the seismic BLRMs for the BS Trillium seismometer. It looks like the problem is probably in the anti-aliasing board. There's some heavy stuff sitting on top of it in the rack, so I'll take a look at it later when someone can give me a hand getting it out.

In detail, after verifying that there are signals coming directly out of the seismometer, I tried to inject a signal into the AA board and see it appear in one of the seismometer channels.

  1. I looked specifically at C1:PEM-SEIS_BS_Z_IN1 (Ch9), C1:PEM-SEIS_BS_X_IN1 (Ch7), and C1:PEM-ACC_MC2_Y_IN1 (Ch27). All of these channels have between 2000--3000 cts.
  2. I tried injecting a 200 mVpp signal at 1.7862 Hz into each of these channels, but the the output did not change.
  3. All channels have 0 cts when the power to the AA board is off.
  4. I then tried to inject the same signal into the AA board and see it at the output. The setup is shown in the first attachment. The second BNC coming out of the function generator is going to one of the AA board inputs; the 32 pin cable is coming directly from the output. All channels give 4.6 V when when the board is powered on regardless of wheter any signal is being injected.
  5. To verify that the AA board is likely the culprit, I also injected the same signals directly into the ADC. The setup is shown in the second attachment. The 32 pin cable is going directly to the ADC. When injecting the same signals into the appropriate channels the above channels show between 200--300 cts, and 0 cts when no signal is injected.
Attachment 1: AA.jpg
AA.jpg
Attachment 2: ADC.jpg
ADC.jpg
  13754   Sat Apr 14 14:42:09 2018 gautamUpdateALSFibers switched out

It looks like the drift in polarization content in the PSL pickoff is actually much higher than that in the EX pickoff. Note that to prevent the P-pol diode from saturating, I put an ND filter in front of the PD, so the Y axis actually has to be multiplied by 10 to compare power in S and P polarizations. If this drift is because of the input (linear) polarization being poorly matched to one of the fiber's special axes, then we can improve the situation relatively easily. But if the polarization drift is happening as a result of time-varying stress (due to temp. fluctuations, acoustics etc) on the (PM) fiber from the PSL fiber coupler to the BeatMouth, then I think this is a much more challenging problem to solve.

I'll attempt to quantify the contribution (in Hz/rtHz) of beat amplitude RIN to phase tracker output noise, which will tell us how much of a problem this really is and in which frequency bands. In particular, I'm interested in seeing if the excess noise around 100 Hz is because of beat amplitude fluctuations. But on the evidence thus far, I've seen the beat amplitude drift by ~15 dB (over long timescales) on the control room network analyzer, and this drift seems to be dominated by PSL light amplitude fluctuations.

Attachment 1: PSLdrift.png
PSLdrift.png
  13753   Fri Apr 13 17:56:26 2018 gautamUpdateALSFibers switched out

I swapped the EX fiber for the PSL fiber in the polarization monitoring setup. There is a lot more power in this fiber, and one of the PDs was saturated. I should really have put a PBS to cut the power, but I opted for putting an absorptive ND1.0 filter on the PD instead for this test. I want to monitor the stability in this beam and compare it to the EX beam's polarization wandering.

  13752   Fri Apr 13 16:59:12 2018 gautamUpdateALSEX green mode-matching

THIS CALCULATION IS WRONG FOR THE OVERCOUPLED CAV.

Summary:

Mode-matching efficiency of EX green light into the arm cavity is ~70*%, as measured using the visibility. 

Details:

I wanted to get an estimate for the mode-matching of the EX green beam into the arm cavity. I did the following:

  1. Locked arm cavities to IR. Ran dither alignment servos to maximize the transmission of IR on both arms. The X arm dither alignment servo needs some touching up, I can achieve higher TRX by hand than by running the dither.
  2. Aligned green PZT mirrors so as to maximize GTRX. Achieved level as 0.47.
  3. Went to EX table and tweaked the two available mode-matching lens positions on their translational stages. Saw a quadratic maximum of GTRX about some equilibrium position (where the lenses are now).
  4. Measured average value of the green PDH reflection DC level whiel green TEM00 mode was locked. P_{\mathrm{locked}} = 716 \mathrm{cts}.
  5. Misaligned ITMX macroscopically. Measured the average value of the green PDH reflection DC level again. P_{\mathrm{misaligned}} = 3800 \mathrm{cts}.
  6. Closed EX Green shutter. Measured the average value of the green PDH reflection DC level. P_{\mathrm{dark}} = 30 \mathrm{cts}.
  7. Modulation depth of the EX PDH was determined to be 90mrad. Based on this, power in sideband is negligible compared to power in the carrier, so I didn't bother correcting for sideband power in reflection.
  8. Mode-matching efficiency calculated as \frac{P_{\mathrm{misaligned}} - P_{\mathrm{locked}}}{P_{\mathrm{misaligned}} + P_{\mathrm{locked}} - 2P_{\mathrm{dark}} }.

Comments:

This amount of mode-matching is rather disappointing - using a la mode, the calculated mode-matching efficiency is nearly 100%, but 70% is a far cry from this. The fact that I can't improve this number by either tweaking the steering or by moving the MM lenses around suggests that the estimate of the target arm mode is probably incorrect (the non-gaussianity of the input beam itself is not quantified yet, but I don't believe this input beam can account for 30% mismatch). For the Y-arm, the green REFL DC level is actually higher when locked than when ITMY is misaligned. WTF?? surpriseOnly explanation I can think of is that the PD is saturated when green is unlocked - but why does the ADC saturate at ~3000cts and not 32000?


This data is almost certainly bogus as the AA box at 1X9 is powered by +/-5VDC and not +/-15VDC. I didn't check but I believe the situation is the same at the Y-end.

3000 cts is ~1V into the ADC. I am going to change the supply voltage to this box (which also reads in ETMX OSEMS) to +/-15V so that we can use the full range of the ADC.


gautam Apr 26 630pm: I re-did the measurement by directly monitoring the REFLDC on a scope, and the situation is not much better. I calculate a MM of 70% into the arm. This is sensitive to the lens positions - while I was working on the EX fiber coupling, I had bumped the lens mounted on a translational stage on the EX table lightly, and I had to move that lens around today in order to recover the GTRX level of 0.5 that I am used to seeing (with arm aligned to maximize IR transmission). So there is definitely room for optimization here.


 

  13751   Fri Apr 13 11:02:41 2018 gautamUpdateALSEX fiber polarization drift

Attachment #1 shows the drift of the polarization content of the light from EX entering the BeatMouth. Seems rather large (~10%). I'm going to tweak the X end fiber coupling setup a bit to see if this can be improved. This performance is also a good benchmark to compare the PSL IR light polarization drift. I am going to ask Steve to order Thorlabs K6XS, which has a locking screw for the rotational DoF. With this feature, and by installing some HWPs at the input coupling point, we can ensure that we are coupling light into one of the special axes in a much more deterministic way. 

Attachment 1: EX_pol_drift.png
EX_pol_drift.png
  13750   Fri Apr 13 00:20:46 2018 ranaUpdatePEMMEDM setup

changed the setpoint of the EX Seismomter T ctrl servo from 35 to 39 C to see if this helps the stability by decreasing the cooldown time constant.

  13749   Thu Apr 12 18:12:49 2018 gautamUpdateALSNPRO channels hijacked

Summary:

  1. Today, the measured IR ALS noise for the X arm was dramatically improved. The main change was that I improved the alignment of the PSL pickoff beam into its fiber coupler.
  2. The noise level was non-stationary, leading me to suspect power modulation of the RF beat amplitude.
  3. I am now measuring the stability of the power in the two polarizations coming from EX table to the PSL table, using the PSL diagnostic connector channels.
  4. The EX beam is S-polarized when it is coupled into the fiber. The PSL beam is P-polarized. However, it looks like I have coupled light along orthogonal axes into the fiber, such that when the EX light gets to the PSL table, most of it is in the P-polarization, as judged by my PER measurement setup (i.e. the alignment keys at the PSL table and at the EX table are orthogonal). So it still seems like there is something to be gained by trying to improve the PER a bit more.

Details:

Today, I decided to check the power coupled into the PSL fiber for the BeatMouth. Surprisingly, it was only 200uW, while I had ~3.15mW going into it in January. Presumably some alignment drifting happened. So I re-aligned the beam into the fiber using the steering mirror immediately before the fiber coupler. I managed to get ~2.9mW in without much effort, and I figured this is sufficient for a first pass, so I didn't try too much more. I then tried making an ALS beat spectrum measurement (arm locked to IMC length using POX, green following the arm using end PDH servo). Surprisingly, the noise performannce was almost as good as the reference! See Attachment #1, in which the red curve is an IR beat (while all others are green beats). The Y arm green beat performance isn't stellar, but one problem at a time. Moreover, the kind of coherence structure between the arm error signal and the ALS beat signal that I reported here was totally absent today.

Upon further investigation, I found that the noise level was actually breathing quite significantly on timescales of minutes. While I was able to successfully keep the TEM00 mode of the PSL beam resonant inside the arm cavity by using the ALS beat frequency as an error signal and MC2 as a frequency actuator, the DC stability was very poor and TRX was wandering around by 50%. So my new hypothesis is that the excess ALS noise is because of one or more of

  • Beam jitter at coupling point into fiber.
  • Polarization drift of the IR beams.

While I did some work in trying to align the PSL IR pickoff into the fiber along the fast (P-pol) axis, I haven't done anything for the X end pickoff beam. So perhaps the fluctuations in the EX IR power is causing beatnote amplitude fluctuations. In the delay line + phase tracker frequency discriminator, I think RF beatnote amplitude fluctuations can couple into phase noise linearly. For such an apparently important noise source, I can't remeber ever including it in any of the ALS noise budgets.

Before Ph237 today, I decided to use my polarization monitoring setup and check the "RIN" of power in the two polarizations coming out of the fiber on the PSL table. For this purpose, I decided to hijack the Acromag channels used for the PSL diagnostics connector Attachment #2 shows that there is fluctuations at the level of ~10% in the p-polarization. This is the "desired" polarization in that I aligned the PSL beam into the fiber to maximize the power in this polarization. So assuming the power fluctuations in the PSL beam are negligible, this translates to sqrt(10) ~3% fluctuation in the RF beat amplitude. This is at best a conservative estimate, as in reality, there is probably more AM because of the non PM fibers inside the beatmouth.

All of this still doesn't explain the coherence between the measured ALS noise and the arm error signal at 100s of Hz (which presumably can only happen via frequency noise in the PSL).

Another "mystery" - yesterday, while I was working on recovering the Y arm green beat signal on the PSL table, I eventually got a beat signal that was ~20mVpp into 50ohms, which is approximately the same as what I measured when the Y arm ALS performance was "nominal", more than a year ago. But while viewing the Y arm beats (green and IR) simultaneously on an o'scope, I wasn't able to keep both signals synchronised while triggering on one (even though the IR beat frequency was half the green beat frequency). This means there is a huge amount of relative phase noise between the green and IR beats. What (if anything) does this mean? The differential noise between these two beats should be (i) phase noise at the fiber coupler/ inside the fiber itself, and (ii) scatter noise in the green light transmitted through the cavity. Is it "expected" that the relative phase noise between these two signals is so large that we can't view both of them on a common trigger signal on an o'scope? surpriseAlso - the green mode-matching into the Y arm is abysmal.

Anyways - I'm going to try and tweak the PER and mode-matching into the X end fiber a little and monitor the polarization stability (nothing too invasive for now, eventually, I want to install the new fiber couplers I acquired but for now I'll only change alignment into and rotation of the fiber coupler on the EX table). It would also be interesting to compare my "optimized" PSL drift to the unoptimized EX power drift. So the PSL diagnostic channels will not show any actual PSL diagnostic information until I plug it back in. But I suspect that the EPICS record names and physical channel wiring are wrong anyways - I hooked up my two photodiode signals into what I would believe is the "Diode 1 Power" and "Laser crystal temperature" monitors (as per the schematic), but the signals actually show up for me in "Diode 2 Power" (p-pol) and "Didoe 1 Temperature" (s-pol).

Annoyingly, there is no wiring diagram - on my todo list i guess...

@Steve - could you please take a photo of the EX table and update the wiki? I think the photo we have is a bit dated, the fiber coupler and transmon PDs aren't in it...

Attachment 1: IR_ALS_20180412.pdf
IR_ALS_20180412.pdf
Attachment 2: BeatMouthDrift.png
BeatMouthDrift.png
Attachment 3: ETMX_20180416.jpg
ETMX_20180416.jpg
  13748   Thu Apr 12 10:15:33 2018 KiraUpdatePEMMEDM setup

Another update. I've changed the on/off button so that it's visible which state it's in. I did that by changing the type of C1:PEM-SEIS-EX_TEMP_SLOWLOOP from ai to bi (I checked the FSS script and copied the entry for the slowloop). Previously, MEDM was giving me an error that it wasn't an ENUM value when I wanted to use a choice button to indicate the value of slowloop, and this solved the issue. I've also added a StripTool button.

Attachment 1: MEDM_3.png
MEDM_3.png
  13747   Wed Apr 11 10:47:26 2018 SteveUpdateSUSsatellite amps labeled

Satellite amplifiers labeled with date. Old labels left on.

Attachment 1: DSC00912.JPG
DSC00912.JPG
  13746   Wed Apr 11 01:34:31 2018 gautamUpdateIOOActivities today

[kevin, gautam]

activities done today - details/plots/data/evidence tomorrow.

  1. Checked XARM loop shape. Updated NB code to fetch POX data from NDS and undo loop shape rather than using calibration filter bank.
  2. Checked POX loop calibration (m/ct). Number I was using was 8e-13. Tonight we measured 9e-13. Updated filter bank.
  3. Tried to get Y arm green ALS going.
    • Improved GTRY from ~0.05 to 0.3.
    • Tried to improve mode matching onto BBPD on PSL table to see a green beat.
    • But we were unsuccessful.
    • I think I got the near and far field alignment right, and the EY laser temp is set such that I can see an IR beat @~30MHz (so green beat should be at 60 MHz).
    • But I couldn't see anything with scope or with HP spec analyzer.
    • More tomorrow. Motivation to get green ALS working is to get some more confidence that the excess noise is indeed on the PSL light.
  13745   Tue Apr 10 15:42:08 2018 KiraUpdatePEMMEDM setup

An update to the screen. I changed the min/max values for some of the parameters, as well as changing the script so that I could specify the integral gain in terms of 1e-5. I've also added this screen to the PEM tab in the sitemap.

Attachment 1: MEDM_2.png
MEDM_2.png
  13744   Tue Apr 10 14:28:44 2018 gautamUpdateIOOFurther debugging

I am working on IMC electronics. IMC is misaligned until further notice.

  13743   Mon Apr 9 23:49:50 2018 ranaUpdatePSLPSL chans

I think we can scrap the 126MOPA channels since they're associated with the Lightwave NPRO and MOPA. We should add the channels that we need for monitoring the Innolight NPRO from the d-sub connector on its controller.

  13742   Mon Apr 9 23:28:49 2018 johannesConfigurationDAQc1psl channel list

I made a list of all the physical c1psl channels to get a better idea for how many acromags we need to replace it eventually. There  3123 unit is the one whose failure had prevented c1psl from booting, which is why it was unplugged (elog post 12852), and its channels have been inactive since. Are the 126MOPA channels used for the current mephisto? 126 tells me it's for an old lightwave laser, but I was checking a few and found that they have non-zero, changing values, so they may have been rewired.

It also hosts some virtual channels for the ISS with root C1:PSL-ISS_ defined in iss.db and dc.db, the PSL particle counter with root C1:PEM- defined in PCount.db  and a whole lot of PSL status channels defined in pslstatus.db. Transfering these virtual channels to a different machine is almost trivial, but the serial readout of the particle counter would have to find a new home.

Long story short - we need:

Function Type # Channels #Channels (no MOPA) # Units # Units (no MOPA)
ADC XT1221 34 21 5 3
DAC XT1541 17 14 3 2
BIO XT1111 19 10 2 1

 



3113 - ADC

C1:PSL-126MOPA_126PWR
C1:PSL-126MOPA_DTMP
C1:PSL-126MOPA_LTMP
C1:PSL-126MOPA_DMON
C1:PSL-126MOPA_LMON
C1:PSL-126MOPA_CURMON
C1:PSL-126MOPA_DTEC
C1:PSL-126MOPA_LTEC
C1:PSL-126MOPA_CURMON2
C1:PSL-126MOPA_HTEMP
C1:PSL-126MOPA_HTEMPSET
C1:PSL-FSS_RFPDDC
C1:PSL-FSS_LODET
C1:PSL-FSS_FAST
C1:PSL-FSS_PCDRIVE
C1:PSL-FSS_MODET
C1:PSL-FSS_VCODETPWR
C1:PSL-FSS_TIDALOUT
C1:PSL-PMC_RFPDDC
C1:PSL-PMC_LODET
C1:PSL-PMC_PZT
C1:PSL-PMC_MODET


3123 - ADC (failed)

C1:PSL-126MOPA_AMPMON
C1:PSL-126MOPA_126MON
C1:PSL-FSS_RCTRANSPD
C1:PSL-FSS_MINCOMEAS
C1:PSL-FSS_RMTEMP
C1:PSL-FSS_RCTEMP
C1:PSL-FSS_MIXERM
C1:PSL-FSS_SLOWM
C1:PSL-FSS_TIDALINPUT
C1:PSL-PMC_PMCTRANSPD
C1:PSL-PMC_PMCERR
C1:PSL-PPKTP_TEMP


4116 - DAC

C1:PSL-126MOPA_126CURADJ
C1:PSL-126MOPA_DCAMP
C1:PSL-126MOPA_DCAMP-
C1:PSL-FSS_INOFFSET
C1:PSL-FSS_MGAIN
C1:PSL-FSS_FASTGAIN
C1:PSL-FSS_PHCON
C1:PSL-FSS_RFADJ
C1:PSL-FSS_SLOWDC
C1:PSL-FSS_VCOMODLEVEL
C1:PSL-FSS_TIDAL
C1:PSL-FSS_TIDALSET
C1:PSL-PMC_GAIN
C1:PSL-PMC_INOFFSET
C1:PSL-PMC_PHCON
C1:PSL-PMC_RFADJ
C1:PSL-PMC_RAMP


XVME-210 - Binary Input

C1:PSL-126MOPA_FAULT
C1:PSL-126MOPA_INTERLOCK
C1:PSL-126MOPA_SHUTTER
C1:PSL-126MOPA_126LASE
C1:PSL-126MOPA_AMPON


XVME-220 - Binary Output

C1:PSL-126MOPA_126NE
C1:PSL-126MOPA_126STANDBY
C1:PSL-126MOPA_SHUTOPENEX
C1:PSL-126MOPA_STANDBY
C1:PSL-FSS_SW1
C1:PSL-FSS_SW2
C1:PSL-FSS_FASTSWEEP
C1:PSL-FSS_PHFLIP
C1:PSL-FSS_VCOTESTSW
C1:PSL-FSS_VCOWIDESW
C1:PSL-PMC_SW1
C1:PSL-PMC_SW2
C1:PSL-PMC_PHFLIP
C1:PSL-PMC_BLANK

  13741   Mon Apr 9 18:46:03 2018 gautamUpdateIOOFurther debugging
  1. I analyzed the data from the free swinging MC test conducted over the weekend. Attachment #1 shows the spectra. Color scheme is same for all panels.
    • I am suspicious of MC3: why does the LR coil see almost no Yaw motion?
    • The "equilibrium" values of all the sensor signals (at the IN1 of the coil input filters) are within 20% of each other (for MC3, but also MC1 and MC2).
    • The position resonance is also sensed more by the side coil than by the LR coil.
    • To rule out satellite box shenanigans, I just switched the SRM and MC3 satellite boxes. But coherence between frequency noise as sensed by the arms remain.
  2. I decided to clean up my IMC nosie budget a bit more.
    • Attachment #2 shows the NB as of today. I'll choose a better color palette for the next update.
    • "Seismic" trace is estimated using the 40m gwinc file - the MC2 stack is probably different from the others and so it's contribution is probably more, but I think this will suffice for a first estimate.
    • "RAM" trace is measured at the CM board input, with MC2 misaligned.
    • The unaccounted noise is evident from above ~8 Hz.
    • More noises will be added as they are measured.
    • I am going to spend some time working on modeling the CM board noise and TF in LTspice. I tried getting a measurement of the transfer function fron IN1 to the FAST output of the CM board with the SR785 (motivation being to add the contribution of the input referred CM board noise to the NB plot), but I suspect I screwed up something w.r.t. the excitation amplitude, as I am getting a totally nonsensical shape, which also seems to depend on my input excitation amplitude. I don't think the output is saturated (viewed during measurement on a scope), but perhaps there are some subtle effects going on.
Attachment 1: MC_Freeswinging.pdf
MC_Freeswinging.pdf
Attachment 2: IMC_NB_20180409.pdf
IMC_NB_20180409.pdf
  13740   Mon Apr 9 16:30:21 2018 KiraUpdatePEMMEDM setup

I created an MEDM screen for the PID control. In addition, I added a new EPICS channel for the setpoint so that it could be adjusted using the MEDM screen.

Edit: forgot to mention the channel name is C1:PEM-SEIS_EX_TEMP_SETPOINT

Edit #2: the path for the MEDM is /opt/rtcds/caltech/c1/medm/c1pem/C1PEM_SEIS_EX_TCTRL.adl

Attachment 1: MEDM_screen.png
MEDM_screen.png
  13739   Mon Apr 9 08:39:39 2018 SteveUpdatePEMM5.3 eq Souther CA

Earth quake M5.3    2018-04-05 19:29:16UTC          Santa Cruz Island, CA

 

Attachment 1: M5.3_Santa_Cruz_Is.CA.png
M5.3_Santa_Cruz_Is.CA.png
Attachment 2: after_M5.3.png
after_M5.3.png
Attachment 3: M5.3vac.png
M5.3vac.png
  13738   Fri Apr 6 22:23:53 2018 KevinUpdateIOOCoil driver noise

 

Quote:

Why is MC2 LR so different from the others???

The previous measurements were made from the coil driver output monitors. To investigate why the MC2 LR coil has less noise than the other coils, I also measured the noise at the output to the coils.

The circuit diagram for the coil driver board is given in D010001 and a picture of the rack is on the 40m wiki here. The coil driver boards are in the upper left quadrant of the rack. The input to the board is the column of LEMOs which are the "Coil Test In" inputs on the schematic. The output monitors are the row of LEMOs to the right of the input LEMOs and are the "FP Coil Volt Mon" outputs on the schematic. The output to the coils "Coil Out" in the schematic are carried through a DB15 connector.

The attachment shows the voltage noise for the MC2 LR coil as well as the UL coil which is similar to all of the other coils measured in the previous measurement. While the LR coil is less noisy than the UL coil as measured at the output monitor, they have the same noise spectrum as measured at the output to the coils themselves. So there must be something wrong with the buffer circuit for the MC2 LR voltage monitor, but the output to the coils themselves is the same as for the other coils.

Attachment 1: MC2_coil_driver.pdf
MC2_coil_driver.pdf
  13737   Fri Apr 6 21:39:09 2018 gautamUpdateIOOMC2 suspension health checkup

While Kevin is working on the MC2 electronics chain - we disconnected the output to the optic (DB15 connector on coil driver board). I decided to look at the 'free' freeswinging MC2 OSEM shadow sensor data. Attachment #1 suggests that the suspension eigenmodes are showing up in the shadow sensors, but the 0.8Hz peak seems rather small, especially compared to the result presented in this elog.

Maybe I'll kick all 3 MC optics tonight and let them ringdown overnight, may not be a bad idea to checkup on the health of the MC suspensions/satellite boxes... [MC suspensions were kicked @1207113574]. PSL shutter will remain closed overnight...

Quote:

Why is MC2 LR so different from the others???

 

Attachment 1: MC2_Freeswinging.pdf
MC2_Freeswinging.pdf
  13736   Fri Apr 6 18:28:57 2018 ranaUpdatePEMPID tuning

Made some changes:

  1. Set P and D gains to zero. We only need slow drift control.
  2. Changed names of the python script and .ini file to distinguish it from the FSS stuff. Lives in scripts/PEM/
  3. removed debug flag from argParse. To run in non-debug mode you use the "-O" option of python as usual.
  4. Fixed the upper/lower limit convention for the heater. Was backwards.
  5. Removed the "rail" function that was defined. We can just use numpy.clip since that's already built in.

There is also now a StripTool file in the scripts/PEM directory which has appropriate channel names and scales for PID loop tuning. Use this file!

I'm leaving it running over the weekend with K_I = -0.003. There is a StripTool on rossa which you can watch. The code itself is running on a tmux session on megatron. Let's ONLY run this code there until we're satisfied that things are good.

Update Sun Apr 8 00:40:11 2018: Lowered gain by factors of 3 down to -0.0001 Saturday afternoon. Seems like still oscillating a bit, now with a ~4 hour period. Setting it to -3e-5 now. Usually we have a linear feedback loop, but our actuation voltage actually gets squared (P = I^2 R) before being integrated to produce temperature. Wonder if we should think of linearizing the feedback control signal to make the loop act nicer.

Update Sun Apr 8 21:09:48 2018: Set K_I = -1e-5 earlier today. Seems to have stabilized nearly, but temperature swings are still +/- 1 K. Will need to add some proportional feedback (K_P) to increase the loop bandwidth, but system is at least sort of stable now. Probably should start construction of EY,BS systems now.

Attachment 1: HeaterTest.png
HeaterTest.png
  13735   Fri Apr 6 16:17:20 2018 KiraUpdatePEMPID tuning

I have been trying to tune the PID and have managed to descrease the oscillations without saturating the actuator. I'm going to model the system to calculate the exact values of P, I and D in order to get rid of the oscillations altogether. I was going to record the data using Data Viewer, but there seems to be some issue with that, so I'm using StripTool for now.

Attachment 1: PID_tuning_progress.png
PID_tuning_progress.png
ELOG V3.1.3-